Kostelnik Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on total fixed manufacturing overhead cost of $562,400, variable manufacturing overhead of $2.30 per machine-hour, and 76,000 machine-hours. The company has provided the following data concerning Job A496 which was recently completed:

Number of units in the job 20
Total machine-hours 80
Direct materials $ 780
Direct labor cost $ 1,560

The unit product cost for Job A496 is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • $29.25

  • $38.95

  • $117.00

  • $155.80
Answer #1

Unit product cost is computed as:

Working note:

1.Predetermined overhead rate per machine hour is computed as:

Predetermined overhead rate=Fixed overhead+variable overhead/machine hours

Fixed overhead=$562,400

Variable overhead=2.30*76000=$174,800

Predetermined overhead rate=$562,400+$174,800/76,000

Predetermined overhead rate=$737200/76,000

Predetermined overhead rate=9.7

2. unit product cost=$3116/20=$155.80